Based on the the tables below.

The public holds 50% of money one (M1) in the form of currency, and the required reserve ratio is 20%.

1a. Calculate the Money Multiplier for question for table 1a
1a. Based on its computed change in money supply what does this imply about the publics desire for holding currency and the money multiplier?

1b. Calculate the Money Multiplier for question for table 1b
1b. Based on its computed change in money supply what does this imply about the publics desire for holding currency and the money multiplier?

Which scenario will contribute more to increase in money?

TABLE 1a

table 1b

Loan proceeds held as Loan proceeds deposited Required Deposits Round Excess reserves Loans currency reserves $500.00 $100.00 $400.00 $400.00 $200.00 $200.00 1 $40.00 2 $200.00 $160.00 $160.00 $80.00 $80.00 $80.00 $16.00 $64.00 $64.00 $32.00 $32.00 $32.00 $6.40 $25.60 $25.60 $12.80 $12.80 4 $12.80 $2.56 $10.24 $10.24 $5.12 $5.12 5 $4.10 $5.12 $1.02 $4.10 $2.05 $2.05 $2.05 $0.41 $0.66 $0.66 $0.33 $0.82 7 $0.82 $0.16 $0.66 $0.66 $0.33 $0.33 8 $0.33 $0.07 $0.26 $0.26 $0.13 $0.13 9 $0.13 $0.03 $0.10 $0.10 $0.05 $0.05 10 $833.25 $166.65 $666.60 $666.60 $333.30 $333.30 Totals
Loan proceeds held as Required Loan proceeds deposited Round Excess reserves Loans Deposits currency reserves $400.00 $500.00 $100.00 $400.00 0.00 $400.00 1 $400.00 $80.00 $320.00 $320.00 $320.00 2 0.00 $320.00 $64.00 $256.00 $256.00 $256.00 3 0.00 $256.00 $51.20 $204.80 $204.80 $204.80 0.00 4 $204.80 $40.96 $163.84 $163.84 0.00 $163.84 5 $163.84 $32.77 $131.07 $131.07 $131.07 0.00 $131.07 $26.21 $104.89 $104.89 $104.86 7 0.00 $104.89 $83.89 $89.89 $20.97 0.00 $83.89 $83.89 $16.78 $67.11 $67.11 $67.11 0.00 $67.11 $13.42 $53.69 $53.69 $53.69 10 0.00 0.00 $446.31 $1785.26 $1785.26 $2231.57 $1785.26 0.00 Totals
Answer #1

In Table 1a: Total new Loans deposited = $333.30 and initial deposit was $500

Hence a deposit of $500 is creating money of $333.30

Hence Money multiplier = 333.3/500 = 0.67

It shows that the higher desire of public to hold money reduces the potential of money multiplier to create money

In Table 1b: Total new Loans deposited = $1785.26 and initial deposit was $500

Hence a deposit of $500 is creating money of $1785.26

Hence Money multiplier = 1785.26/500 = 3.57

It shows that the when there is no desire of public to hold money it increases the potential of money multiplier to create money

It can be said that scenario in Table 1b where public desire to hold money is zero will contribute more to create money supply.